Find a general form of an equation of the line that has y-intercept 3/2 and passes through the center of the circle with endpoints of a diameter A(4, -3) and B(-2, 7).


*I tried to solve this by doing point-slope form since I had a point or two, and the slope already given to me. When I did that I ended up with 2x+3y=17 somehow, which I know isn't right. Then I tried finding the midpoint, and calculating slope to put into point slope form, which also didn't work. I know that the answer is x-2y=-3, but I would like to know the steps and approach to get to the right answer. Thank you!

Answer by Alan3354(69443)

Find a general form of an equation of the line that has y-intercept 3/2 and passes through the center of the circle with endpoints of a diameter A(4, -3) and B(-2, 7).


*I tried to solve this by doing point-slope form since I had a point or two, and the slope already given to me. When I did that I ended up with 2x+3y=17 somehow, which I know isn't right. Then I tried finding the midpoint, and calculating slope to put into point slope form, which also didn't work. I know that the answer is x-2y=-3, but I would like to know the steps and approach to get to the right answer.
===========
Find the center of the circle.
C is (1,2)
------
Find a general form of an equation of the line that has y-intercept 3/2 and passes through the center of the circle with endpoints of a diameter A(4,-3) and B(-2,7).
-----
Find the slope, m
m = diffy/diffx = (1.5 - 2)/(0 - 1)
m = 1/2
----
y = mx + b
2 = (1/2)1 + b
b = 3/2
----
--> y = (1/2)x + 3/2
x - 2y = -3